40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 226     Entry time: Mon Jan 7 09:01:39 2008
Author: steve 
Type: Update 
Category: SUS 
Subject: BS sus damping restored 
The BS sus damping was lost at 8am Sunday morning.
Attachment 1: bssdl.jpg  57 kB  | Hide | Hide all
bssdl.jpg
ELOG V3.1.3-